I need help with finance homework!!!! Thank you so much!!!!!
What is the present value of the following future amounts?
A.$800 to be received 10 years from now discounted back to the present at 10 percent.
B.$300 to be received 5 years from now discounted back to the present at 5 percent

C.$1,000 to be received 8 years from now discounted back to the present at 3 percent

D.$1,000 to be received 8 years from now discounted back to the present at 20 percent

Answer by rfer(16322)

P=FV/(1+r)^t
P=800/1+.1)^10
P=800/2.59374246
P=$308.43
Try the others yourself.
Good luck.